### v2

# GenAINet: Enabling Wireless Collective Intelligence via Knowledge Transfer and Reasoning

Abstract
[…]
GenAI and communication networks are expected to have groundbreaking synergies in 6G. Connecting generative artificial intelligence (GenAI) agents over a wireless network can potentially unleash the power of collective intelligence and pave the way for artificial general intelligence (AGI). However, current wireless networks are designed as a “data pipe” and are not suited to accommodate and leverage the power of GenAI. In this paper, we propose the GenAINet framework in which distributed GenAI agents communicate knowledge (high-level concepts or abstracts) to accomplish arbitrary tasks. We first provide a network architecture integrating GenAI capabilities to manage both network protocols and applications. Building on this, we investigate effective communication and reasoning problems by proposing a semantic-native GenAINet. Specifically, GenAI agents extract semantic concepts from multi-modal raw data, build a knowledgebase representing their semantic relations, which is retrieved by GenAI models for planning and reasoning. Under this paradigm, an agent can learn fast from other agents’ experience for making better decisions with efficient communications. Furthermore, we conduct two case studies where in wireless device query, we show that extracting and transferring knowledge can improve query accuracy with reduced communication; and in wireless power control, we show that distributed agents can improve decisions via collaborative reasoning. Finally, we address that developing a hierarchical semantic level Telecom world model is a key path towards a network of collective intelligence.
I Introduction
[…]
To deploy LLMs on resource limited devices, various techniques have been proposed, such as model compression, inference acceleration, and parameter-efficient fine tuning. For instance, the T5-770M model has been shown to outperform the PaLM-540B in some tasks [3]. Moreover, paged attention proposed in vLLM can reduce the average inference time of LLMs down to 1/24 compared to vanilla methods [4]. Furthermore, QLoRA [5] is an efficient fine-tuning technique which can quantize an LLM to a 4-bit version with a small set of additional learnable low-rank adapters tuned using quantized weights. Although efficient in model size and inference time reductions, such techniques can degrade LLM’s performance, e.g., in terms of generalization capability, due to the non-modular nature of LLMs.
To use LLMs in complex tasks solving, a multi-agent framework is required. An agent can observe the environment, plan a sequence of actions, create high-level reflections of experience in a memory stream, and formulate plans. For example, BabyAGI is a task-driven autonomous agent framework that can generate, execute and prioritize tasks in real-time. Auto-GPT can chain together LLM “thoughts” in an infinite loop of reasoning [6]. On the other hand, multi-agent systems are shown to largely enhance LLMs’ capability in task solving [7]. For example, CAMEL explores two role-playing communicative agents with autonomous cooperation [8]. In generative agents, believable human behaviors are simulated in a sandbox environment, where multiple agents interact in natural language to complete spontaneous tasks [9]. However, communication, computing, and storage efficiency are not considered in these frameworks, which are critical factors in wireless networks.
Motivated by the reasons above, in this paper, we propose the GenAINet framework in which multiple GenAI agents communicate knowledge to perform effective reasoning and solve arbitrary tasks of particular applications, e.g., networks of vehicles, smart grid, and internet of things. In GenAINet, agents utilize LLMs to learn a high-level concept or abstract of data, plan reasoning path, and execute it to achieve a goal. The agents communicate knowledge to help each other’s planning and decision. In doing so, the performance can be improved with less communication and computing costs. The main contributions of this article are as follows:
  • A GenAINet architecture is discussed and possible approaches to integrate LLMs are proposed. We propose a unified agent architecture to manage both network protocols and applications;
  • A communication paradigm for GenAINet, built on semantic concepts and multi-path reasoning, is proposed. With semantic concepts, multi-modal raw data can be embedded on a common semantic space. With multi-step planning, we explain how an LLM can execute and optimize reasoning paths towards effective decisions. Semantic concepts learned from past reasoning and decision can be sent to or retrieved by remote agents, to improve independent or collaborative planning;
  • We investigate two case studies of wireless GenAINets. In the first case of question answering on a mobile device, we show transferring knowledge from a teacher to a student agent can improve query accuracy and communication efficiency; in the second case of wireless power control, we show that distributed GenAI agents can collaborate with reasoning to find a power allocation solution for a targeted data rate;
  • We discuss challenges and explore future research directions on GenAINet, including building a hierarchical world model with multi-level abstractions and planning to ground the model in real network scenario.
II GenAI Network and Agent Architectures
Wireless GenAI agents can emulate human-like decision making process, providing a path towards sophisticated and adaptive GenAI networks. It can bring autonomy to network protocols and network applications. To achieve this, we propose prospect architectures for GenAI networks and GenAI agents, as illustrated in Fig. 1.
Figure 1: Proposed wireless GenAI network and agent architecture with protocol and application management.
II-A Wireless GenAI Network Architecture
In a wireless network, a GenAI agent can potentially act as an autonomous controller in the following scenarios.
Network protocols: GenAI agents can orchestrate network resources and control network functions in radio access, core, and transport domains. It can interface with network protocols to control power, resource, traffic, etc. The GenAI network protocol architecture can include:
  • Hierarchical architecture: GenAI agents can be deployed in multi-level network controllers, sending policies to the data network. For instance, a commander agent operates in the RAN intelligent controller (RIC), while multiple executor agents operate in the network elements (NE). The commander deconstruct intents into subtasks and assigns to the executors.
  • Distributed architecture: GenAI agents can be co-located with distributed NEs and manageme locally the data and control plans. The interaction between agents is self-organized, where agents exchange past decisions or knowledge to collaboratively accomplish the tasks.
In autonomous networks, GenAI agents can break down high-level intents, plan actionable tasks, generate network control strategies, and refine them from feedback.
Network applications: GenAI agents can bring autonomy to network applications, such as autonomous vehicles and robots. Wireless networks allow GenAI agents to collaboratively perform various tasks, such as remote sensing, control, and planning. The GenAI network application architecture can include:
  • Independent architecture: Each application is controlled by individual GenAI agents. For example, in a vehicular network, the network agents control base stations, user equipments (UEs) and the traffic agents control cars, traffic lights.
  • Converged architecture: GenAI agents can jointly control network protocols and applications. For example, a car agent can jointly control engine, steering and communication protocols with other cars, such that multiple cars can collaboratively improve the traffic flow.
The hierarchical and independent architecture can be built over existing 5G wireless networks where applications are over-the-top services. On the other hand, the distributed and converged architectures are more disruptive for 6G, fully implementing the protocol and application convergence with multi-agent network.
II-B Wireless GenAI Agent Architecture
To enable LLMs to interact with real-world scenarios, an agent has to be of a distributed structure, decision-wise and information-wise. An autonomous agent generally comprises four components: perception, action, planning, and memory. In what follows, we explain how these components can be designed to implement the proposed vision.
Perception is the part of an agent gathering useful and relevant information to an LLM. Possible information sources for perception include: 1) information sensed from the environment, e.g., channel state information (CSI), locations of user equipment (UE) and BS in a network, or on an autonomous car, the traffic density, speed and visual data of streets; 2) information received from other agents, such as messages reporting their state or responses to some requested information, or strategies learned from past decisions under different states. The information can be represented by using different modalities, from raw text, images, to abstracted graphs. The perceptor should fuse and encode raw data in a common embedding space, for the LLM to perform further decision-making.
Action is the component managing all available tools of an agent. Actions generated by the LLM should be adapted to the target functions or interfaces to execute. For a wireless GenAI agent, possible actions may include: 1) information generated to complete a task, such as the responses to a query from user, or messages sent to another agent to complete a task; 2) executing predefined interfaces and toolkits, such as the function call in a MAC scheduler, configurations in the radio amplifier, or commands to adjust the car speed and steering. Furthermore, the action in a multi-agent network includes actions taken by observing the environment and communication actions (messages) to be sent to other agents. Both should be generated from a learned joint strategy.
Planning is the process of creating and optimizing actions over time to achieve a goal including: 1) sub-task decomposition, i.e., breaking down a high-level goal into sub-goals and actionable tasks; 2) synchronize and prioritize existing tasks between agents; 3) self-reflection, which evaluates and criticizes the past decisions to optimize the policy. Planning allows LLMs to solve complex tasks via multi-step reasoning, such as chain-of-thoughts (CoTs). Furthermore, such planning and reasoning can be conducted collaboratively among multiple agents by leveraging computing resources and tools on different devices efficiently.
Memory is the module for storing short-term experience and long-term knowledge for LLMs’ future planning and decisions. Short-term experience is the selected history of observations, actions, thoughts, and conclusions of agents. It is necessary for agents to efficiently reuse or adapt existing solutions or mechanisms instead of reasoning and extracting from scratch. Moreover, the long-term knowledge contains facts, methods, and contents which could be general or domain-specific depending on the role of an agent. Having such knowledge could effectively reduce the chance of an LLM to produce non-reliable or non-factual responses by using retrieval-augmented techniques. Possible knowledge representations and how LLMs leverage them to enhance the quality of inference will be covered in next section. Empowered by LLMs’ knowledge, GenAI agents can release the need for environment modeling or problem formulation, while finding the reasoning path towards task completion. This makes GenAI effective in handling complex, unseen problems.
III Wireless GenAINet with Knowledge driven Communication and Reasoning
Existing LLMs and GenAI agents operate on raw data. For example, agents memorize and exchange information in natural language. Communication between AI models on the network edge is typically done by exchanging model weights, gradients, or hyper-parameters. However, such design schemes become inefficient when massive GenAI agents are connected. First, raw data contains a large amount of redundant information which is inefficient for communication; Second, optimizing LLMs over networks for different tasks is energy-consuming and may induce large latency. To address the aforementioned issues, we propose a semantic-native GenAINet. Empowered by LLM’s abstraction and planning capabilities, the agents communicate semantic concepts or abstracts learned from reasoning and decision making. In doing so, agents can learn to communicate or collaborate to achieve a goal, with reduced communication, computing and storage costs.
III-A Semantic knowledge representation
LLMs have been trained to understand semantics. Extracting knowledge from raw data to a concept space allows LLMs to effectively compose information for different tasks, which leads to better generalization. Knowledge includes information abstracted from a wide range of experiences. In GenAINet, this can be retrieved from common sources such as the web. Besides knowledge can be learned from an agent’s plan-and-execute experiences. For example, a car agent may learn to decelerate when observing other cars merging from the side lane. Another example, is when a solver agent may share with other solver agents an intermediate result about its calculation or computation process. Both retrieved and learned knowledge are thus vital for effective decision-making or generation.
Wireless GenAINet being constrained by communication, computing, and storage resources, knowledge transferred and stored in the network should satisfy minimality and sufficiency, that a minimal amount of information can solve a wide range of tasks or problems effectively. To realize this, we can decompose knowledge into two components: a) common knowledge in a domain which can be injected into LLMs via fine-tuning; b) evolving knowledge operating on top of common knowledge for specific tasks, which should be transferred in the network and used by LLMs. In GenAINet, managing common and evolving knowledge properly is the key to achieve high efficiency. In what follows, three possible representations for knowledge which are relevant for communication between LLMs are discussed.
[…]
KG is a structured representation of relations between real objects or abstracted concepts. These relations can describe connections and causalities between entities in natural language, making it easily accessible for an LLM. Two vertices would be connected by an edge if they are correlated and the weight of the edge may indicate the degree of correlation. knowledge graph (KG) can assist an LLM to perform fast information retrieval, and efficient reasoning based on the structure of entities. Moreover, KG can also help LLMs to understand and generate more contextually relevant and coherent responses for complex tasks requiring factual responses.
[…]
TE: Topological models, e.g., hypergraphs, simplicial complexes and cell complexes, can represent the intrinsic structure of data. They can model the low-order (vertices) and high-order (simplices, cells) relations between syntactic objects or abstracted concepts. Unlike conventional neural networks learning data structure on 1D sequence or 2D grid, topological embedding (TE) can exploit more flexible and dynamic structure. Moreover, the high-order topological structure can represent more complex and causal relations, such as different word combinations by semantic features. Therefore, TE can exploit more implicit latent structures than KG. The geometric features allows for flexible structure composition according to the changes of semantic features, which saves communication and computing resources.
III-B Multi-modal semantic reasoning
Semantic knowledge allows GenAI agents to compress, transfer, and retrieve new information. However, conventional structures built on the observation space lack generalization in new scenarios and domains. Therefore, building structures on a semantic concept space is essential. This requires aligning multi-modality raw data on a common embedding space and extracting semantic concepts. Techniques such as contrastive learning and multi-modal cross-attention (e.g., ImageBind [10]) have been used to align multi-modal data, extendable to RF signals as well. After training cross-modality encoders, semantic concepts can be extracted from raw data uniformly.
Semantic knoweldgebase can be built by learning a topological structure connecting semantic concepts. The first step is to exploit hidden semantic concepts from observations. Connections between semantic concepts can be learned from their semantic relations and causalities, which can be represented by different topological structures (e.g., edges and cliques). With logical combination of explicit concepts we can discover hidden concepts. This process can learn all possible semantic structures from raw data, then serving as a “world model” in the cloud. Subsequently, minimal and sufficient semantic concepts can be retrieved from semantic knoweldgebase to serve specific tasks, e.g., by clustering the concepts according to the semantic information required for a task.
The semantic concepts are then used on agents to perform planning and reasoning, for making decisions or generating contents. CoT prompting could be used to decomposes a problem into coherent sequences serving as intermediate steps, with an interpretable window suggesting how it might reach the final solution, and gives the agent opportunity to debug when a step goes wrong [11]. CoT could be generalized to sequential or non-sequential complex structures such as tree or graph of thoughts, with rewards to optimize the path. Furthermore, the decomposed sub-tasks of a complex task can be assigned to different agents and be solved separately. Finally the outputs from each sub-task can guide agents to complete the task jointly.
Fig. 2 illustrates the process of multi-modal semantic reasoning. Semantic concepts $c$ are learned from multi-modal raw data. A semantic knowledgebase (KB) representing the semantic relations between concepts on topological structure is built. When an agent receives an input, it plans a tree of thought states $T$. In each state it retrieves concepts from semantic KB and produce an action. The final output is generated after completing the planned states. The agent can observe rewards at each state, optimize the reasoning path and the semantic concept structures in the KB.
Figure 2: Proposed pipeline of multi-modal semantic extraction, retrieval and reasoning on GenAINet agents.
III-C Semantic-native GenAINet
With semantic knowledge and semantic reasoning, GenAI agents can collectively learn communication protocols and decision strategies for solving different tasks, leading to an efficient GenAINet. GenAI agents can be used for various tasks on connected devices, from generating content to making decisions. An effective communication protocol should be part of the reasoning process, guiding agents to achieve a goal, such as the accuracy of responses, the accumulated reward of actions, and so on. Furthermore, an agent should effectively manage the knowledge to transmit, memorize, and update its model according to the information freshness and energy cost.
GenAINet can be implemented in a teacher-student or a distributed paradigm. The teacher-student paradigm is applicable for communication between cloud, edge, and device agents, where different sizes of LLMs are deployed. Specifically, a teacher agent with LLM trained on universal knowledge guides a student agent to perform specific tasks via knowledge transfer. To reduce communication cost, the teacher agent extracts knowledge in specific domains and transfer to the student, which is efficient for devices performing specific tasks (e.g., routing, traffic control). The distributed paradigm can be used for communication between mobile devices, machines, vehicles, where agents have similar capabilities and need to collaboratively complete tasks. Each agent has a memory of emergent knowledge which is learned from its past experience. During planning, the agent can retrieve or communicate knowledge with others to optimize decisions and update local knowledge. In a long term, the agents will share a common KB and make decisions locally, where communication costs can be minimized until observing new scenarios or tasks. Compared to standard communication systems, GenAINet is resource efficient and task effective. For example, in a remote query scenario the teacher-student paradigm can reduce load and latency in sending responses; in an autonomous vehicle (AV) scenario the distributed paradigm can improve driving safety and communication reliability.
Fig. 3 illustrates an example of GenAINet communication between AVs. The car agent utilizes LLMs as knowledge retriever (sender) and decision maker (actor). Communication is part of knowledge retrieval or sharing with a remote agent, which is planned according to actions and observations in a reasoning path. This makes the information exchange efficient and effective for decision making. The retriever aggregates information from local and remote KB, pass it to the actor to generate driving commands. The actor optimizes reasoning path from environment feedback. For example, it can reward the CoT to find the best path and action. Once complete, the agent extracts knowledge (e.g., learned rules) from the recent planning and store it in the memory.
We shown here three typical use cases of semantic communication between car agents. The first case is semantic compression from multi-modal sensory data. LLMs can produce semantic latent on source textual data for lossless generation [12]. With multi-modal LLMs, semantic compression can be used to gather information efficiently from remote sensor with different field-of-views or resolutions, to improve downstream tasks such as accident detection, navigation. The second case is collaobrative knowledge, where agents exchange the abstracts and concepts from memory. This can be achieved by remote retrieval, where the agent performs semantic similarity search on the request from remote agent, and send the related information. The third case is semantic reasoning, where agents find the consequence of past actions and exchange with others. This can reduce agents’ reasoning latency and improve decision reliability. Since long-term knowledge is more generalized than short-term experience, the communication cost can be reduced.
Figure 3: Three use case of semantic native GenAINet: compression, control, and reasoning.
IV Case Studies of GenAINet
In this section, we show how the GenAINet can be applied to two examples namely, wireless device query and wireless power control. Specifically, we want to show: 1) How GenAI agents can efficiently transfer knowledge to perform on-device query; 2) How GenAI agents can collaborate with reasoning to solve a wireless network problem.
IV-A Knowledge transfer for wireless device query
Question answering (QnA) is a typical application of LLMs. Deploying LLMs close to the end user can largely reduce the latency and traffic burden from massive connections. Despite that light-weight LLMs are built with efficient inference techniques, they exhibit poor performance in specialized domains compared to regular LLMs. RAGs can enhance LLM’s knowledge with external database while introducing storage cost and higher latency. Envisioned from GenAINet, we propose a semantic knowledge driven on-device query. As shown in Fig. 4, In the cloud agent, its LLM extracts the related context for the queries from cloud data and compress them into latent representations to build a semantic KB. It is then sent to the LLM on a device to answer questions using RAG. The device can feedback questions to improve semantic extraction.
We have performed experiments on the TeleQnA [13], including 10k Telecom domain QnA from research and standard materials. We extract the questions’ sources and build a large vector DB in the cloud. An initial batch of sampled questions are fed into GPT-3.5 to perform retrieval from the DB, under instruction prompts to generate contexts related to the questions, and build the semantic KB for RAG on device. Table I shows that the semantic KB significantly improves the TeleQnA accuracy in all categories compared to the base Llama-7B, by reducing 27% exchanged information compared to cloud based QnA. The use case shows that an LLM can extract semantic knowledge and assist other LLMs to complete a task which is more efficient than sending raw data.
Figure 4: Cloud LLMs extract semantic knowledge from raw text and send to LLM on devices for QnA with RAG.
QnA category GPT-3.5-turbo accuracy Semantic RAG accuracy Reduced bits exchange
Lexicon 96 100 31.09
Research overview 66.35 72.12 26.61
Research publication 66.98 83.72 27.43
Standard overview 64.52 85.48 26.59
Standard specification 56.38 71.28 23.91
Overall 66 80 27.13
TABLE I: QnA accuracy and reduced bits exchange (%), compared to exchanging raw QnA data.
IV-B Collaborative reasoning for wireless power control
Conventional power control paradigm resorts to an optimization problem by modeling the environment and finding the best power level leading to assigned performance metric (rate, energy). With the proposed paradigm, the solving is performed by exploiting LLMs’ knowledge of wireless network (acquired during pre-training), to reason over the path from observed rate to best power. This means in particular that with GenAI agents we may release online training.
We consider a scenario of 20 paired Tx-Rx users randomly placed in a $100m^{2}$ area. Users share a common spectrum where interference is encountered. Each user pair is associated with an LLM agent, including an LLM with plan-and-solve prompting, plus a memory of past data rate observations and power allocation actions. We aim to reduce the data rate gaps towards a target so as to minimize the total power in the network. The radio environment information is unknown to the agents and we instruct LLMs to exploit its memory and its telecom knowledge. In each iteration each LLM should decide its power and explain its decision.
The mean absolute error of data rate of all users towards the target is shown in Fig. 5. We evaluated 3 scenarios: 1) standalone agents without interactions; 2) agents share memory of past observations and actions; 3) agents share an explanation of the past decisions. It can be seen that the standalone agents find the power optimization path towards the targeted rate after several rounds. With memory sharing, agents learn slowly to understand the propagation and interference environment and adjust their power in the right direction. Finally, the explanation from other agents in previous rounds can help LLMs to find better power allocation scheme quickly. The example shows that sharing reasoning outputs can effectively help LLMs to improve decision making without online training.
Figure 5: Performances of data rate gap towards target with example prompts to LLM agents on power control.
V Challenges and Opportunities
We envision GenAINet as a vital part for enabling collective intelligence in 6G. However, there are several challenges which opens up new research opportunities.
LLMs are auto-regressive generative models pre-trained on natural text to predict the next token. Therefore LLMs hardly predict high-level semantics, which introduces high information redundancy and computation cost in training and inference. Besides, LLMs’ generalization capability is usually limited in domains not included in the training data. RAG and fine-tuning are commonly used to enhance LLMs’ knowledge while being expensive and inflexible. Furthermore, grounding LLMs’ knowledge to real world representation is challenging. Reinforcement learning is studied to optimize LLM online, which is not suitable for computational limited devices.
World model is a hierarchical, modular model predicting future representations of the state of the world [14]. It is trained to predict the high-level abstract instead of raw data. For example, I-JEPA [15] is trained to predict missing embedding on an image, which shows less training data and computational effort. Furthermore, hierarchical JEPA (H-JEPA) allows to learn higher level abstract representation which is effective for long-term prediction and eliminate the irrelevant details. Built on H-JEPA, hierarchical planning can be done by predicting state transition on abstract space, which could handle uncertain environment with minimal cost. JEPA framework gives promises for deploying world model on wireless devices than LLMs. Since RF propagation is more abstract than text, training RF-JEPA could be more efficient than RF-GPT. However, JEPA on multi-modality and planning still require further research.
While GenAI agents are used in various domains, applying them in wireless networks is still challenging. First, a network has a complex hierarchical architecture from RF to service layers, making LLMs difficult to decompose tasks for every network element. Second, orchestrating numerous agents’ behavior is difficult for networks deployed in large scale geographical areas. Therefore, current LLMs are used only in small domains, mostly service layer. Moreover, future networks require much higher reliability and robustness, where the uncertainty of GenAI models is problematic. Finally embedding RF signals into LLMs is challenging, primarily due to the unavailability of appropriate large datasets and the inherent nature of RF signals, which is both spectral and spatial. This fundamentally differs from the textual data.
VI Conclusion
In this paper, we introduced the GenAINet, a knowledge driven communication and reasoning network as a promising enabler for collective intelligence. The architecture of GenAINet leverages unified LLM-powered GenAI agents to optimize both network protocols and applications. Our semantic-native framework can learn structure of semantic concept or abstract from multi-modal raw data, and build semantic knowledgebase to achieve effective communication and reasoning. Agents utilize LLMs with minimal semantic concepts retrieved locally and remotely through semantic communications to effectively plan reasoning paths and produce effective decisions for a task. We investigated a use case of transferring semantic knowledge between teacher and student agents to improve on-device query with less communication costs, as well as an example of distributed agents communicate with reasoning to improve wireless power control. We demonstrated that our multi-agent GenAINet can unleash the power of collective intelligence, and addressed that developing a semantic-native Telecom world model is an essential path towards network collective intelligence.
Biographies
Hang Zou (hang.zou@tii.ae) is a Researcher at Technology Innovation Institute, UAE.
Qiyang Zhao (qiyang.zhao@tii.ae) is a Principal Researcher at Technology Innovation Institute, UAE.
Lina Bariah (lina.bariah@ieee.org) is an Adjunct Professor at Khalifa University, UAE.
Yu Tian (yu.tian@tii.ae) is a Researcher at Technology Innovation Institute, UAE.
Mehdi Bennis (mehdi.bennis@oulu.fi) is a Professor at University of Oulu, Finland.
Samson Lasaulce (samson.lasaulce@ku.ac.ae) is a Chief Research Scientist with Khalifa University (UAE) and a CNRS Director of Research (France).
Mérouane Debbah (merouane.debbah@ku.ac.ae) is a Professor at Khalifa University and a Senior Advisor at Technology Innovation Institute, UAE.
Faouzi Bader (carlos-faouzi.bader@tii.ae) is the Senior Director of Telecom Unit at Technology Innovation Institute, UAE.